# Revisiting and Maximizing Temporal Knowledge in Semi-supervised Semantic Segmentation

### v1

@@ 4.4 Discussion
PrevMatch 93 84 61 87 71 68 93 84 89 21 89 57 84 86 77 82 51 82 46 84 58
$\Delta$ 2 0 2 -2 0 0 1 4 1 13 1 2 -1 1 2 2 -3 0 13 4 -6
TABLE XI: Class-wise IoU scores for Cityscapes with a 1/8 label partition using a ResNet-101 encoder.
@@
PrevMatch 98 84 92 60 63 63 71 80 92 63 95 82 64 95 83 88 81 69 77
$\Delta$ 0 2 0 4 3 1 0 1 0 3 0 0 1 0 0 1 2 1 0
4.4.1 Analysis of Performance and Training Stability
Class-wise IoU Scores and Qualitative Evaluation. Tables X and XI list the category-wise IoU scores. In particular, Table X on Pascal VOC shows that the proposed method achieves notable performance gains for the chair and sofa classes, which were particularly challenging for the UniMatch baseline. In addition, Table XI on the Cityscapes dataset shows that the proposed method achieves the largest performance improvements for the wall, fence, and terrain classes, which are the lowest performing among the 19 classes in UniMatch. In addition to the quantitative results, the qualitative results shown in Fig. 3 corroborate these findings, revealing consistent improvements in the same categories. Thus, these results suggest that utilizing previous knowledge helps prevent the catastrophic forgetting problem described in Section 3.2, even in semi-supervised semantic segmentation scenarios. We further investigate this problem in the subsequent analysis.
Training Stability of Poorly Behaved Classes. To further explore the catastrophic forgetting problem in poorly behaved classes, we visualize changes in terms of pseudo-label accuracy and validation IoU scores during training, as depicted in Fig. 4. In the first row (chair class), the training curve of the baseline pseudo-label accuracy exhibits significant fluctuations, particularly showing a sudden sharp performance drop at approximately 50 epochs. Although the pseudo-label accuracy recovers slightly thereafter, the validation score does not. In the second row (sofa class), the training curve of the baseline exhibits more severe fluctuations and sharper and more drastic performance drops than in the first row. In contrast, the proposed method shows a smoother training curve without significant fluctuations in either category. This indicates that the proposed training procedure aids in achieving stable optimization for poorly behaved classes that suffer from the forgetting problem.
Figure 4: Training curves for the chair and sofa classes, illustrating variations in pseudo-label pixel accuracy and validation IoU scores. The experiment is conducted on the 92-label partition of Pascal VOC.
Training Stability Across Different Label Partitions. Fig. 5 illustrates the effect of the proposed method on the changes in the validation scores throughout the training process. In fewer label settings (92 and 183), the baseline (blue) exhibits significant fluctuations in terms of performance compared to the proposed method (orange). Moreover, when considering the epoch that achieves the best performance, the baseline method struggles to converge consistently across epochs and tends to become trapped in local minima prematurely. This issue is particularly pronounced in scenarios with fewer labels. In contrast, the proposed method consistently converges across epochs without significant fluctuations. Finally, the consistent outperformance of our method over the baseline across almost all training epochs in label partitions suggests that the positive effects of previous guidance outweigh any negative effects. Note that the positive and negative effects refer to the cases (3) and (2), respectively, described in the last paragraph of Section 3.2.
The results presented in this subsection suggest the following implications: (1) Training with fewer labels leads to instability in pseudo-label predictions, particularly for poorly behaved classes. (2) The proposed training procedure aids in stabilizing pseudo-label predictions while mitigating the forgetting problem in poorly behaved classes. (3) In other words, leveraging previous knowledge helps alleviate the issues of tight coupling and catastrophic forgetting, which can hinder stable learning in semi-supervised segmentation settings.
@@ 4.4.1 Analysis of Performance and Training Stability
TABLE XII: Comparison of time spent using existing methods with a ResNet-101 encoder. Training time per epoch and GPU memory usage were measured using the same environment (two A6000 GPUs) and hyperparameters, such as batch size, in a 92-label partition. We used the open-sourced code provided by the authors.
Method #Trainable Networks Training Time (m) GPU Memory Usage (G) Pascal
92 183 366
UniMatch 1 7.8 21.9 75.2 77.2 78.8
Diverse Co-T. 3 22.1 40.5 75.4 76.8 79.6
PS-MT 1 24.9 28.6 65.8 69.6 76.6
ours ($K$=1) 1 8.4 22.7 76.4 77.5 79.3
ours ($K$=3) 1 9.1 22.7 77.0 78.5 79.6
4.4.2 Efficiency Evaluation
To investigate the complexities of the proposed method, we measure the training time per epoch and GPU memory usage. In Table XII, Diverse Co-training, based on the co-training approach, slightly outperforms UniMatch, which operates on a single trainable network, in terms of performance. However, Diverse Co-training requires approximately three times more training time and twice as much memory, resulting in limited scalability. Moreover, PS-MT, a dual Mean Teacher-based method, requires training time comparable to that of Diverse Co-training due to its complex components, despite being a single trainable network. In contrast, the proposed method significantly outperforms existing methods, with a slight cost increase. In particular, PrevMatch ($K$=3) outperforms UniMatch by 1.8% in terms of mIoU (92 label setting) while requiring only a slight increase in complexities in training time and GPU usage, suggesting that PrevMatch can efficiently provide diverse pseudo-labels. Therefore, owing to its computational efficiency and the simplicity of its pipeline reusing the same inputs, the proposed method can be easily integrated into any semi-supervised learning method.
5 Conclusion and Future Work
In this work, we introduced the PrevMatch framework, which leverages temporal knowledge obtained during training to efficiently address the issues of tight coupling and confirmation bias that impede stable semi-supervised learning. The main contributions of PrevMatch include revisiting the use of temporal knowledge and maximizing its effectiveness. Specifically, we directly utilized previous models to provide additional pseudo-label guidance, referred to as previous guidance, to the student network. In addition, we developed a highly randomized ensemble strategy that enhances the reliability and diversity of the previous guidance while minimizing the increase in computational complexity. Experiments were conducted on four benchmark semantic segmentation datasets, revealing that PrevMatch significantly outperforms existing state-of-the-art methods across different evaluation protocols. Furthermore, our findings indicate that leveraging temporal knowledge facilitates stable optimization, particularly for classes that exhibit poor performance and fluctuations. Finally, the computational efficiency and compatibility of the proposed method facilitate its seamless integration into recent semi-supervised semantic segmentation methods.
@@ 5 Conclusion and Future Work
In future work, we will investigate the capability of PrevMatch to address the domain adaptation problem. This problem poses a more challenging self-training task due to domain discrepancies between the labeled and unlabeled images, a commonly encountered problem in real-world applications. In addition, we observed a phenomenon where significant fluctuations in pseudo-label accuracy for several classes negatively affect generalization ability. For instance, pseudo-label accuracy recovers slightly after a sharp drop in performance; however, the validation score does not. Although the proposed method has shown that it can mitigate these issues, a more in-depth investigation is required for scenarios involving many classes and imbalanced distributions, such as the COCO and ADE20K datasets. Therefore, we intend to explore this phenomenon extensively across different classes, in terms of the relationship between training instability and generalization ability.
Acknowledgments
This research was supported by Brain Korea 21 FOUR. This research was also supported by Korea University Grant (K2403371) and Korea TechnoComplex Foundation Grant (R2112653).

### v2

@@ 8 Additional Analysis
+ PrevMatch 93 84 61 87 71 68 93 84 89 21 89 57 84 86 77 82 51 82 46 84 58
[…]
Table 15: Class-wise IoU scores for Pascal VOC using a ResNet-50 encoder.
@@
+ PrevMatch 98 84 92 60 63 63 71 80 92 63 95 82 64 95 83 88 81 69 77
[…]
Table 16: Class-wise IoU scores for Cityscapes using a ResNet-101 encoder.
8.1 Class-wise IoU Scores and Qualitative Evaluation
Tabs. 15 and 16 list all category-wise IoU scores. In particular, Tab. 15 on Pascal VOC shows that the proposed method achieves notable performance gains for the chair and sofa classes, which were particularly challenging for the UniMatch baseline. In addition, Tab. 16 on the Cityscapes dataset shows that the proposed method achieves the largest performance improvements for the wall, fence, and terrain classes, which are the lowest performing among the 19 classes in UniMatch. In addition to the quantitative results, the qualitative results shown in Fig. 3 corroborate these findings, revealing consistent improvements in the same categories. Thus, these results suggest that utilizing previous knowledge helps prevent the catastrophic forgetting problem, even in semi-supervised semantic segmentation scenarios.
@@ 8.1 Class-wise IoU Scores and Qualitative Evaluation
8.2 Training Stability Across Different Label Partitions
Fig. 4 illustrates the effect of the proposed method on the changes in the validation scores throughout the training process. In fewer label settings (92 and 183), the UniMatch baseline (blue) exhibits significant fluctuations in terms of performance compared to the proposed method (orange). Moreover, when considering the epoch that achieves the best performance, the baseline method struggles to converge consistently across epochs and tends to become trapped in local minima prematurely. This issue is particularly pronounced in scenarios with fewer labels. In contrast, the proposed method consistently converges across epochs without significant fluctuations.
Meanwhile, as mentioned in Sec. 3.2 of the main paper, given the standard and previous guidance, four scenarios can be considered based on whether they are correct or incorrect (standard-previous): (1) correct-correct, (2) correct-incorrect, (3) incorrect-correct, and (4) incorrect-incorrect. Through case (3), the network receives an additional opportunity to be guided in the right direction, away from the wrong one. However, model training may also be hindered through case (2), leading to significant fluctuations. Nevertheless, as shown in Fig. 4, the consistent outperformance of our method over the baseline across almost all training epochs in different label partitions suggests that the positive effects (i.e., case (3)) of previous guidance outweigh any negative effects (i.e., case (2)).
9 Limitations and Future Work
@@
We showed that the proposed method is effective on several benchmark datasets that share the domain between labeled and unlabeled images. However, it has not been investigated in domain adaptation problems that involve domain discrepancies between the labeled and unlabeled data. This problem poses a more challenging self-training task, a commonly encountered problem in real-world applications. In future work, we will investigate the capability of temporal knowledge in the domain adaptation problem. In addition, we observed a phenomenon where significant fluctuations in pseudo-label accuracy for several classes negatively affect generalization ability. For instance, pseudo-label accuracy recovers slightly after a sharp drop in performance; however, the validation score does not. Although the proposed method has shown that it can mitigate these issues, a more in-depth investigation is required for scenarios involving many classes and imbalanced distributions, such as the COCO and ADE20K datasets. Therefore, we intend to explore this phenomenon extensively across different classes, in terms of the relationship between training instability and generalization ability.